find the x and y intercepts and graph the equation plotting the intercepts. please show all of your work.
-7x-8y=56


Answer by nanobot101(2)

-----
to solve for y, let x be zero.
-7(0)-8y=56.
-8y=56
y=-7
to solve for x, let y be zero.
-7x-8(0)=56
-7x=56
x=-8.
now just draw a graph, and plot the points (-8) on the x-axis and (-7) on the y-axis. make a line, and there you go.